OpenStudy (aimmebgood):

1/x - 1/y over 1/x + 1/y simplify?

OpenStudy (anonymous):

Do the numerator and denominator separately. Get a common denominator for the numerator that is xy. Do same for denominator.

OpenStudy (anonymous):

1/x - 1/y = y/xy - x/xy Similar "xy" for the denominator. Then, the "xy" denominator in the numerator and the denominator will cancel. Show your work so far.

OpenStudy (anonymous):

-1/(x y)+1/x+1/y.

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

No, that's not it. Aimee are you working on this? How far are you getting?

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ x }-\frac{ 1 }{ y } }{ \frac{ 1 }{ x }+\frac{ 1 }{ y } } = \frac{ \frac{ y }{ xy }-\frac{ x }{ xy } }{ \frac{ y }{ xy }+\frac{ x }{ xy } } = \frac{ \frac{ y-x }{ xy } }{ \frac{ y+x }{ xy } } = \frac{ y-x }{ y+x }\]